# Tug-of-war between idiom’s figurative and literal meanings in LLMs

### v3

@@ Abstract
Idioms present a unique challenge for language models due to their non-compositional figurative meanings, which often strongly diverge from the idiom’s literal interpretation. This duality requires a model to learn representing and deciding between the two meanings to interpret an idiom in a figurative sense, or literally. In this paper, we employ tools from mechanistic interpretability to trace how a large pretrained causal transformer (LLama3.2-1B-base) deals with this ambiguity. We localize three steps of idiom processing: First, the idiom’s figurative meaning is retrieved in early attention and MLP sublayers. We identify specific attention heads which boost the figurative meaning of the idiom while suppressing the idiom’s literal interpretation. The model subsequently represents the figurative representation through an intermediate path. Meanwhile, a parallel bypass route forwards literal interpretation, ensuring that a both reading remain available. Overall, our findings provide a mechanistic evidence for idiom comprehension in an autoregressive transformer. [^1] Code and dataset are available at https://github.com/sori424/idiom_processing under the MIT License.
$\dagger$\$$\dagger$\$footnotetext: Co-corresponding authors.
@@ 1 Introduction
Figure 1: The figurative and literal interpretations are highlighted in the block and path. We find three main steps for idiom processing: 1 Idiom retrieval step: Early layers (i.e., 0-3) attention and MLP are actively retrieving the idiom’s figurative meaning while storing both figurative and literal interpretations in the residual stream. 2 Selective interpretation step: At the token immediately following the idiom span, the model begins to encode a representation that favors the figurative interpretation over the literal one, starting from the middle layers. 3 Interpretation routing: For final prediction, the model passes literal interpretation via both a direct compositional semantic path ( 3-2), as well as the intermediate pathway that prioritizes the figurative meaning ( 3-1 figurative path).
@@
Idioms challenge standard semantic composition because they are, by definition, multi-word sequences whose figurative meanings diverge from the composition of their literal parts Beck and Weber (2020). For instance, “kick the bucket” conveys the figurative meaning “to die”, yet its literal sense “physically kicking a container” remains semantically available. This ambiguity often leads large language models (LLMs) to misinterpret idioms, favoring their literal sense over the intended figurative one Kabra et al. (2023); Liu et al. (2023); Knietaite et al. (2024). Psycholinguistics has been investigated how humans understand idioms, whether they access the idiomatic meaning directly or first have to access the literal meaning of the idiom and only in a later stage suppress that meaning and access the figurative meaning. These findings have lead to various models describing human idiom processing Gibbs (1980); Cacciari and Glucksberg (1991); Bobrow and Bell (1973); Cacciari and Tabossi (1988).
@@
Yet for LLMs, it’s unclear how they process idioms, potentially limiting further improvements in their idiom comprehension. One explanation may lie in how the model transforms raw input token embeddings into richer representations Tian et al. (2023); Haviv et al. (2022); Dankers et al. (2022). Within the autoregressive transformer, token representations are incrementally refined from the initial input embeddings to the final representation used for next token prediction Vig et al. (2020). This transformation is achieved through a series of residual layers, which consist of multi-head self-attention (MHSA) and MLP sublayers, each adding their outputs to update the representation. This multi-stage transformation suggests that idiom processing might follow a multi-step transformation process. Haviv et al. (2022) demonstrated that the early layers of transformer models store and retrieve idiomatic information, while the higher layers enhance model confidence. Meanwhile, Dankers et al. (2022) showed that during machine translation of idioms, the transformer model recognizes idiomatic expressions and processes them more as single lexical units. This involves grouping the idioms together using attention sublayers and reducing their interaction with surrounding context. However, both studies lack a detailed mapping of the specific model components and causal pathways underlying these idiom processing behaviors.
@@
In this paper, we first aim to identify how the idiom is interpreted by the transformer (i.e., LLama3.2-1B-base), where its figurative meaning is located, and how it is retrieved by the model. To this end, we pinpoint specific components of the model that have specialized in processing idioms by boosting their figurative meaning, while suppressing the compositional semantic literal interpretation. Having identified these, we proceed to investigating the information flow in the model, with the goal of identifying specific pathways through which the figurative meaning is passed forward to the final prediction, and whether this differs from the pathways of passing forward information about an idiom’s literal interpretation.
@@
We address these questions through knockout analyses Nanda et al. (2023); Wang et al. (2022); Geva et al. (2023), where we separately knock out activations of each component (layer-wise MHSA and MLP, individual attention heads) in the model to observe their importance for retrieving idiom’s figurative meaning. Furthermore, to trace the flow of these interpretations, we employ activation patching experiments Wang et al. (2022); Meng et al. (2022); Hanna et al. (2023); Conmy et al. (2023); Stolfo et al. (2023), in which we ablate one or both interpretations from specific pathways within the model.
Our main findings are summarized in Figure 1: we identify three main steps of idiom processing (1) retrieval of the figurative meaning of the idiom in the early layers, (2) selective encoding of the two interpretations at intermediate layers after the idiom span token position, and (3) two alternative paths for routing the idiom’s meaning to the final prediction, via a direct path for the literal meaning and an intermediate pathway that preferentially amplifies figurative interpretation over the other. In summary, our main contributions are:
  • We propose a next token prediction task that distinguishes figurative vs. literal interpretations using interpretation-specific tokens (Section 2).
  • We show that early layers contribute to retrieve the idiom’s figurative meaning, and that these are significantly different from general semantic processing (Section 4).
  • We identify a selection mechanism at the token immediately following the idiom, where its residual activation shifts toward the figurative interpretation and then propagates through the intermediate layers (Section 5).
  • We also uncover a bypass pathway that channels the literal interpretation directly into the final prediction, circumventing the intermediate route (Section 5.2).
@@ 2 Idiom interpretation task
We are interested in teasing apart the language model’s processing of the literal vs. figurative interpretation of ambiguous idiom expressions. We therefore formalize a task and construct a dataset to investigate the interpretation shift in the model. In order to allow for effective probing, we embed the idiom into a sentence template, constructed such that the interpretation is revealed via next token prediction. The template [^2] X is instantiated with various pronouns (he/she/it/they); the word “would” is inserted only in some of the idioms to make the sentence sound more fluent; idioms are morphologically fit into the sentence; depending on sg/pl form of X, the second part of the template contained was or were. ‘X (would) IDIOM because X was so/too/a/the’ can be applied to all of the selected idioms while preserving the idiom’s ambiguity. It elicits a causal completion, which we expect to differ between the literal and the figurative meaning of the idiom: for the figurative meaning of the idiom “kick the bucket”, we would expect completions that provide causes for the figurative meaning “to die”, e.g., old/sick, while for the literal meaning, we would expect completions that provide reasons for “kicking an object”, e.g. angry/mad. Figure 2 illustrates this process and selected examples can be found in Appendix Table 2.
Figure 2: Pipeline for extracting figurative ($C_{f}$) and literal ($C_{l}$) candidate tokens for an idiomatic sentence.
Since there is no single ground truth continuation that reveals either the figurative or the literal reading, we automatically construct sets of multiple tokens that are consistent with either the literal or the figurative meaning. The tokens in each set are mutually exclusive and selected based on their high likelihood of being associated with either the figurative or literal meaning of the idiom. In the following, we describe each step of the process of generating the dataset.
1. Idiom Extraction: We select 245 idioms from a psycholinguistic paper Cronk et al. (1993), each of which received high literality ratings in a human rating study.
@@
.[$s_{a}$] He would kick the bucket because he was so
@@
.[$s_{l}$] He would kick the container because he was so
3. Token Set Generation: To generate sets of next-word completions that are indicative of the figurative vs. the literal interpretation, we prompt the LLaMa3.3-70B model with disambiguated paraphrases ($s_{f}$, $s_{l}$) and obtain logits from the model with greedy decoding. Let $z^{(f)}$ and $z^{(l)}$ denote the logits corresponding to $s_{f}$ and $s_{l}$ as input, respectively. We then compute the element-wise difference between the two sets of logits as follows:
$\Delta z_{v}=z_{v}^{(f)}-z_{v}^{(l)},$
where $v\in V$ is the vocabulary size of the model (i.e., $V=$128,256). Since tokens that score high in one context tend to score low in the other, $\Delta z_{v}$ captures each token’s context-specific relevance. The 20 tokens that have highest $\Delta z_{k}$ are chosen as the set $C_{f}$ of tokens that indicate a figurative interpretation and the 20 tokens that have lowest $\Delta z_{k}$ values are chosen as the set $C_{l}$ that indicates a literal interpretation.
4. Validation: To make sure that the Llama3.2-1B-base model has prior knowledge of the selected idioms, and the capability of discerning the context of the given sentence, we filter out the generated instances that satisfy the following two conditions:
(1) Given $s_{a}$ as a prompt, the cumulative probabilities of predicting a figurative candidate as a next token must exceed that of the literal candidate. This metric quantifies that the model has the knowledge of the idiom:
$\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{a})>\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{a}).$
(2) Given $s_{f}$ and $s_{l}$ as prompt, the cumulative probabilities of predicting next token candidate matching the context exceeds that of the opposite context:
$\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{f})>\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{f}),$
$\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{l})>\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{l}).$
110 out of the original set of 245 idioms passed these criteria. All following analyses are based on this subset of 110 validated idioms.
@@ 3.1 Information flow in transformers
Given an input token sequence $\mathbf{t}=[t_{1},\dots,t_{T}]$ over a vocabulary $V$, each token is embedded as $\mathrm{x}_{i}^{0}=\mathrm{RoPE}(\mathrm{emb}(t_{i}),i)\in\mathbb{R}^{d}$, where $\mathrm{RoPE}$ applies rotary positional encoding to inject position $i$ Su et al. (2024), and $d$ is the model’s hidden dimension. From layer 0, these vectors $\mathrm{x}_{i}^{0}$ are carried forward and accumulated over the subsequent $L$ layers. At each layer $\ell$, the hidden state for token $i$ is updated from its previous value $\mathrm{x}_{i}^{\ell-1}$ by adding the multi-head self-attention (MHSA) output $a_{i}^{\ell}$ and MLP output $m_{i}^{\ell}$:
@@
MHSA SublayersEach attention layer’s output $\mathrm{a}^{\ell}_{i}$ aggregates information of multiple parallel attention heads $h$. The parameter matrices of each MHSA sublayer comprise three projection matrices $W^{Q},\,W^{K},\,W^{V}\in\mathbb{R}^{d\times d}$, and $W_{O}\in\mathbb{R}^{d\times d}$. The columns of each projection matrix and rows of the outputs matrix can be split into $H$ equal parts, corresponding to the number of attention heads $W_{Q}^{\ell,j},\,W_{K}^{\ell,j},\,W_{V}^{\ell,j}\in\mathbb{R}^{d\times\tfrac{d}{H}}$, and $W_{O}^{\ell,j}\in\mathbb{R}^{\tfrac{d}{H}\times d},\,j=[1,H]$. Let $X^{\ell-1}\in\mathbb{R}^{N\times d}$ be the matrix of token representations at layer $\ell-1$. This allows the MHSA output as a sum of matrices, each induced by single attention head:
$a_{i}^{\ell}=\sum_{j=1}^{H}A^{\ell,j}\bigl{(}X^{\ell-1}\,W_{V}^{\ell,j}\bigr{)}\,W_{O}^{\ell,j}.$
After $L$ layers, the final hidden state $\mathrm{x}_{T}^{L}$ is mapped to vocabulary logits via an unembedding matrix $W_{U}\in\mathbb{R}^{|V|\times d}$, so that
$y=W_{U}\,\mathrm{x}_{T}^{L}\in\mathbb{R}^{|V|}$
and $\mathrm{softmax}(y)$ defines the next-token distribution. Thus we can view a transformer as a computational graph $G\colon\mathcal{X}\to\mathcal{Y}$, where token embeddings, MHSA sublayers, MLP sublayers, and the unembedding all communicate via the residual stream. We refer to Vaswani et al. (2017); Geva et al. (2023) for further details.
@@ 3.3 Activation patching
Activation patching is a technique to identify which activations in a model contribute to a particular output Wang et al. (2022); Meng et al. (2022); Hanna et al. (2023); Conmy et al. (2023); Stolfo et al. (2023). Instead of knocking out a component, we replace it with specific representations. This involves running the model on input A, then selectively replacing (patching) certain activations with those obtained from a paired input B:
@@
3.4 Metric for measuring interpretation
@@
$F(s)\;=\;\sum_{c_{f}\in C_{f}}p(c_{f}\mid s);\;L(s)\;=\;\sum_{c_{l}\in C_{l}}p(c_{l}\mid s),$
where $C_{f}$ and $C_{l}$ are the sets of tokens corresponding to figurative and literal interpretation, respectively. After applying an intervention (knockout or activation patching), we quantify its causal effect:
@@
where $\Delta I<0$ means the intervention disrupts the corresponding interpretation, showing the targeted component is essential for that meaning. In contrast, $\Delta I=0$ indicates no effect and $\Delta I>0$ implies that the intervened component is not required (or may even been inhibited) that particular reading.
4 Localizing idiom meaning retrieval
@@ 4.1 Probing sublayers through knockout
To locate the mechanisms for retrieving figurative meanings from the model’s parameters, we knock out the activations at idiom span tokens by replacing them with their mean values computed over all $s_{a}$. By knocking out the activations across each layer, we assess how the probabilities of predicting specific candidate tokens ($\Delta F(s_{a}),\,\Delta L(s_{a})$) change. To narrow down the points to which we can ascribe a functional role for a specific computation component, we limit the intervention to the MLP and MHSA in each block.
@@
Figure 3: Sublayer-wise interpretation shift $\Delta I(s)$ after ablating activations at idiom span, for contexts $s\in\{s_{a},s_{f},s_{l}\}$. Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$, $\Delta L(s_{l})$, $\Delta F(s_{f})$ with 95% confidence intervals. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant difference between $\Delta F(s_{a})$ and the others (paired $t$-test, $p<0.05$). The difference at * marked layer is larger than the average difference across all layers.
@@
Results Figure 3(a) shows by how much the model’s interpretation of the idiom changes when an MLP sublayer is knocked out. We observe substantial drops of $\Delta F(s_{a})$ when knocking out early MLP sublayers (0-2): from $-0.57$ to $-0.25$. At the same time, we see an increase in the probability of the literal completions: $\Delta L(s_{a})$ rises above zero, from $0.37$ to $0.18$. And these shifts are significantly different from 0-2 layers ($p<0.05$). This indicates that early layer MLPs are diagnostic for reading out the figurative meaning while potentially suppressing the literal one. For later layers ($\geq$ 4), MLP sublayer knockout does not affect the model’s interpretation, which converges to zero.
@@
Figure 3(b) shows the effect of attention knockout, which is similar to what we observed with MLP sublayer knockout. Specifically, knocking out MHSA leads to a drop in $\Delta F(s_{a})$ and increase in $\Delta L(s_{a})$ in the early layers (0-3). The most significant difference between $\Delta F(s_{a})$ and $\Delta L(s_{a})$ is in layer 1, where $\Delta F(s_{a})$ dropped by $-0.36$ and $\Delta L(s_{a})$ increased by $0.43$ ($p<0.05$). After layer 4, knocking out MHSA has little to no effect on interpretation, as the values converge toward zero.
@@
Together, the two interventions track the 1 idiom’s figurative meaning retrieval: Early MHSA layers (0-3) gather and bind token-wise interactions, while early MLP layers (0-2) transform those bound representations into figurative meaning, while suppressing the literal meaning. This early layer idiom meaning retrieval aligns with the findings of Haviv et al. (2022), who argue that idiomatic information is accessed in the initial layers of LLMs during inference. After layer 4, $\Delta F(s_{a})$ and $\Delta L(s_{a})$ stabilize to zero, indicating that later sublayers are not adding new information and that the representations of the idiom span tokens are no longer used by any subsequent tokens.
4.2 Disentangling idiom specific processing from general semantic processing
To validate whether early layers in MHSA and MLP components are specifically essential for idiom processing, or generally important for semantic interpretation of a sentence, we conduct additional knockout experiments on literal and figurative paraphrases ($s_{l},\,s_{f}$). If general compositional semantic processing does not need to rely on the idiom retrieval components, we would expect that knocking out them when processing non-ambiguous sentences (i.e., paraphrases) would not disrupt the inference.
@@ 4.2 Disentangling idiom specific processing from general semantic processing
ResultsFigure 3(c) shows that the drop of knocking out early MLP sublayers is significantly smaller in layers 1 and 2 for the two paraphrases $\mathrm{min}$($\Delta L(s_{l})$, $\Delta F(s_{f})$)$=-0.11$ than for the idiomatic expression, $\mathrm{min}$($\Delta F(s_{a})$)$=-0.36$ ($p<0.05$). While, at layer 0, the knockout effect is large for all expressions, as in this case, the MLP block is processing the semantics of the sentence.
Figure 3(d) shows a similar effect for the early attention layers (0-2) where literal paraphrases are not strongly affected by knockout; $\mathrm{min}$($\Delta L(s_{l})$, $\Delta F(s_{f})$)$=-0.08$, $\mathrm{min}$($\Delta F(s_{a})$)$=-0.36$ ($p<0.05$). Taken together, these results reveal that early MLP and MHSA sublayers are essential for the model’s retrieval of an idiom’s figurative meaning, and that this process is notably different for idiomatic phrases from general semantic processing. Knocking out the early layers preserves the general semantic comprehension, but disrupts its ability to interpret idiom expressions. This dissociation implies that the model activates a specialized component that is necessary for idiom understanding, yet remains inactive during non-ambiguous literal semantic processing.
4.3 Identifying attention heads specific to retrieving figurative meanings

### v4

# Tug-of-war between idioms’ figurative and literal interpretations in LLMs

@@ Abstract
Idioms present a unique challenge for language models due to their non-compositional figurative interpretations, which often strongly diverge from the idiom’s literal interpretation. In this paper, we employ causal tracing to systematically analyze how pretrained causal transformers deal with this ambiguity. We localize three mechanisms: (i) Early sublayers and specific attention heads retrieve an idiom’s figurative interpretation, while suppressing its literal interpretation. (ii) When disambiguating context precedes the idiom, the model leverages it from the earliest layer and later layers refine the interpretation if the context conflicts with the retrieved interpretation. (iii) Then, selective, competing pathways carry both interpretations: an intermediate pathway prioritizes the figurative interpretation and a parallel direct route favors the literal interpretation, ensuring that both readings remain available. Our findings provide mechanistic evidence for idiom comprehension in autoregressive transformers [^1] Code and dataset are available at https://github.com/sori424/idiom_processing under the MIT License..
[…]
@@ 1 Introduction
Figure 1: The figurative and literal interpretations are highlighted in the blocks and paths. We find three main steps for idiom processing: 1 Idiom retrieval step: Early layers (i.e., layers 0-3) attention and MLP are actively retrieving the idiom’s figurative interpretation while storing both figurative and literal interpretations in the residual stream. 2 Selective interpretation step: At the token immediately following the idiom span, the model begins to encode a representation that favors the figurative interpretation over the literal one, starting from the middle layers. 3 Interpretation routing: For final prediction, the model passes literal interpretation via both a direct compositional semantic path ( 3-2), as well as the intermediate pathway that prioritizes the figurative interpretation ( 3-1 figurative path).
@@
Idioms pose a challenge to standard semantic composition because, as multi-word expressions, their figurative meanings don’t follow from the literal meanings Beck and Weber (2020). For instance, “kick the bucket” conveys the figurative interpretation “to die”, yet its literal sense “physically kicking a container” remains semantically available. Psycholinguistics has investigated how humans understand idioms, whether they access the figurative interpretation directly or first have to access the literal interpretation of the idiom and only in a later stage suppress that interpretation and access the figurative interpretation. These findings have led to various models describing human idiom processing Gibbs (1980); Cacciari and Glucksberg (1991); Bobrow and Bell (1973); Cacciari and Tabossi (1988).
@@
Yet for LLMs, it’s unclear how they process idioms, potentially limiting further improvements in their idiom comprehension Kabra et al. (2023); Liu et al. (2024); Knietaite et al. (2024). One explanation may lie in how the model transforms raw input token embeddings into richer representations Tian et al. (2023); Haviv et al. (2023); Dankers et al. (2022). Within the autoregressive transformer, token representations are incrementally refined through a series of residual layers, which consist of multi-head self-attention (MHSA) and MLP sublayers, each adding their outputs to update the representation Vig et al. (2020). This multi-stage transformation suggests that idiom processing proceeds via a sequential shift from compositional to non-compositional representations Haviv et al. (2023); Dankers et al. (2022).
@@
In this paper, we first aim to identify how the idiom is interpreted by a transformer Llama3.2-1B (we also run experiments on Llama3.1-8B, Qwen2.5-0.5B, Qwen2.5-7B), where its figurative interpretation is located, and how it is retrieved by the model compared to its literal counterpart. To this end, we pinpoint specific components of the model that have specialized in processing idioms by boosting their figurative interpretation, while suppressing the compositional semantic literal interpretation. Having identified these, we proceed to investigating the information flow in the model, with the goal of identifying specific pathways through which the figurative interpretation is passed forward to the final prediction, and whether this differs from the pathways of passing forward information about an idiom’s literal interpretation. Additionally, given that the context plays a key role in disambiguating the interpretation of an idiom Mi et al. (2025); Fazly et al. (2009); Dankers et al. (2022); Holsinger (2013), we investigate how context interacts with lexical information of idioms.
@@
We address these questions through knockout analyses Nanda et al. (2023); Wang et al. (2022); Geva et al. (2023), where we separately ablate activations of each component (layer-wise MHSA and MLP, individual attention heads) in the model to observe their importance for retrieving idiom’s figurative interpretation, and contextual disambiguation. Furthermore, to trace the flow of these interpretations, we employ activation patching experiments Wang et al. (2022); Meng et al. (2022); Hanna et al. (2023); Conmy et al. (2023); Stolfo et al. (2023), in which we ablate one or both interpretations from specific pathways within the model. Figure 1 summarizes our main findings. Our contributions are:
  • We find that early layers and idiom specific heads causally retrieve figurative interpretations, while suppressing the corresponding literal counterparts (Section 4).
  • We identify a selection mechanism at the token immediately following the idiom, where an intermediate path carries a strong flow toward the figurative interpretation. In parallel, a bypass pathway preferentially carries literal information directly to the final prediction, circumventing the intermediate pathway (Section 5).
  • We localize where context resolves idiomatic ambiguity, showing that early MLP/MHSA layers suppress figurative retrieval under literal context, while later MHSA integrates surrounding tokens to refine the context consistent interpretation (Section 6).
@@ 2 Idiom interpretation task
We conduct our experiments in two setups. First, we investigate how language models process idiomatic expressions, which permit both figurative and literal interpretations. In the second setup, we introduce disambiguating context by adding preceding sentences.
We formalize a next-token prediction task and construct a dataset. Idioms are embedded into a sentence template designed to reveal their interpretation through continuation ‘X (would) IDIOM because X was so/too/a/the’ [^2] Here, X is instantiated with pronouns and morphologically adapted; X is instantiated with various pronouns (he/she/it/they); the word “would” is inserted only in some of the idioms to make the sentence sound more fluent; idioms are morphologically fit into the sentence; depending on sg/pl form of X, the second part of the template contained was or were.. This template preserves idiom ambiguity while eliciting causal completions (e.g., for ‘kick the bucket’, figurative completions include ‘old/sick’, while literal completions include ‘angry/mad’.) Since there is no single ground truth continuation, we automatically construct token sets consistent with either the figurative or literal interpretation. The tokens in each set are mutually exclusive and chosen for their high likelihood of association with the respective interpretation. Below, we outline the dataset construction process, and see Figure 9:
1. Idiom Extraction: We select 245 idioms from a psycholinguistic paper Cronk et al. (1993), each of which received high literality (humans rated the literal meaning as plausible). This ensures ambiguity, which is necessary to evaluate the model’s change between literal and figurative interpretations.
@@
$s_{a}$] He would kick the bucket because he was so
@@
$s_{l}$] He would kick the container because he was so
3. Token Set Generation: To generate sets of next-word completions that are indicative of the figurative vs. the literal interpretation, we prompt the Llama3.3-70B-instruct model with paraphrases ($s_{f}$, $s_{l}$) and obtain logit distributions. Let $z^{(f)}$ and $z^{(l)}$ denote the logits corresponding to $s_{f}$ and $s_{l}$ as input, then we compute the element-wise difference between the two sets $\Delta z_{v}=z_{v}^{(f)}-z_{v}^{(l)}$, where $v\in V$ is the vocabulary size of the model. Since tokens that score high in one paraphrase tend to score low in the other, $\Delta z_{v}$ captures each token’s interpretation-specific relevance. We take the 20 tokens with the largest $\Delta z_{k}$ as the figurative candidate set $C_{f}$, and the 20 with the smallest $\Delta z_{k}$ as the literal candidate set $C_{l}$.
4. Context Generation: Using each paraphrased sentence, we generate a preceding context sentence that disambiguates the interpretation of the subsequent idiom. The examples below illustrate the generated figurative context ($FC$) and literal context ($LC$).
[…]
.[FCFC] His breath rattled after every step. He would die because he was so
[…]
.[LCLC] The vending machine ate his coins. He would kick the container because he was so
@@ 3.1 Information flow in transformers
Given an input token sequence $\mathbf{t}=[t_{1},\dots,t_{T}]$ over a vocabulary $V$, each token is embedded as $\mathbf{x}_{i}^{0}=\mathrm{PE}(\mathrm{emb}(t_{i}),i)\in\mathbb{R}^{d}$, where $\mathrm{PE}$ denotes a positional encoding function that injects position $i$ into the token embedding Vaswani et al. (2017), and $d$ is the model’s hidden dimension. From layer 0, these vectors $\mathrm{x}_{i}^{0}$ are carried forward and accumulated over the subsequent $L$ layers. At each layer $\ell$, the hidden state for token $i$ is updated from its previous value $\mathrm{x}_{i}^{\ell-1}$ by adding the multi-head self-attention (MHSA) output $a_{i}^{\ell}$ and MLP output $m_{i}^{\ell}$:
@@
MHSA SublayersEach attention layer’s output $a_{i}^{\ell}$ aggregates information of multiple parallel attention heads $h$. The parameter matrices of each MHSA sublayer comprise three projection matrices $W_{Q}^{\ell},\,W_{K}^{\ell},\,W_{V}^{\ell}\in\mathbb{R}^{d\times d}$, and $W_{O}^{\ell}\in\mathbb{R}^{d\times d}$. The columns of each projection matrix and rows of the outputs matrix can be split into $H$ equal parts, corresponding to the number of attention heads $W_{Q}^{\ell,j},\,W_{K}^{\ell,j},\,W_{V}^{\ell,j}\in\mathbb{R}^{d\times\tfrac{d}{H}}$, and $W_{O}^{\ell,j}\in\mathbb{R}^{\tfrac{d}{H}\times d},\,j=[1,H]$. Let $X^{\ell-1}\in\mathbb{R}^{N\times d}$ be the matrix of all token representations at layer $\ell-1$, $A^{\ell,j}\in\mathbb{R}^{N\times N}$ encodes the weights computed by the $j$-th attention head at layer $\ell$, where it’s masked to a lower triangular matrix. The MHSA output is the sum over individual attention heads matrices:
$a_{i}^{\ell}=\sum_{j=1}^{H}\big(A_{i}^{\ell,j}X^{\ell-1}W_{V}^{\ell,j}\big)W_{O}^{\ell,j}.$
After $L$ layers, the final hidden state $\mathrm{x}_{T}^{L}$ is mapped to vocabulary logits via an unembedding matrix $W_{U}\in\mathbb{R}^{|V|\times d}$, so that $y=W_{U}\,\mathrm{x}_{T}^{L}\in\mathbb{R}^{|V|}$, and $\mathrm{softmax}(y)$ defines the next‐token distribution. Thus we can view a transformer as a computational graph $G\colon\mathcal{X}\to\mathcal{Y}$, where token embeddings, MHSA sublayers, MLP sublayers, and the unembedding all communicate via the residual stream. We refer to Vaswani et al. (2017); Geva et al. (2023) for further details.
@@ 3.3 Activation patching
Activation patching is a technique to identify which activations in a model contribute to a particular output Wang et al. (2022); Meng et al. (2022); Hanna et al. (2023); Conmy et al. (2023); Stolfo et al. (2023). This involves running the model on input A, then selectively patching certain activations with those obtained from a paired input B:
@@
3.4 Metric for measuring interpretation shift
@@
$F(s)\;=\;\sum_{c_{f}\in C_{f}}p(c_{f}\mid s);\;L(s)\;=\;\sum_{c_{l}\in C_{l}}p(c_{l}\mid s).$
After applying an intervention, we quantify its causal effect:
@@
where $\Delta I<0$ means the intervention disrupts the corresponding interpretation, implying the component is necessary; $\Delta I=0$ shows no effect; $\Delta I>0$ implies the component is not required (or may even inhibit) that particular reading.
4 Localizing idiom’s figurative interpretation retrieval
We deliberately preserve idiomatic ambiguity by analyzing idioms without the context, so we can localize where, and through which components, the model shifts between figurative and literal interpretations.
@@ 4.1 Probing sublayers through knockout
To locate the mechanisms for retrieving figurative interpretations from the model’s parameters, we knock out the activations at idiom span tokens by replacing them with their mean values computed over all $s_{a}$. By knocking out the activations across each layer, we assess how the probabilities of predicting specific candidate tokens ($\Delta F(s_{a}),\,\Delta L(s_{a})$) change. To narrow down the points to which we can ascribe a functional role for a specific computation component, we limit the intervention to the MLP and MHSA in each block.
@@
Figure 2: Sublayer-wise interpretation shift $\Delta I(s)$ after ablating activations at idiom span, for sentences $s\in\{s_{a},s_{f},s_{l}\}$. Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$, $\Delta L(s_{l})$, $\Delta F(s_{f})$ with 95% confidence intervals. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant difference between $\Delta F(s_{a})$ and the others (paired $t$-test, $p<0.05$). The difference at * marked layer is larger than the average difference across all layers.
@@
Results Figure 2(a) shows by how much the model’s interpretation of the idiom changes when an MLP sublayer is knocked out. We observe substantial drops of $\Delta F(s_{a})$ when knocking out early MLP sublayers (0-2). At the same time, we see an increase in the probability of the literal completions: $\Delta L(s_{a})$ rises above zero. And these shifts are significantly different in 0-3 layers ($p<0.05$). This indicates that early layer MLPs are causally important for reading out the figurative interpretation while potentially suppressing the literal one. For later layers ($\geq$ 4), MLP sublayer knockout does not affect the model’s interpretation, where the corresponding patching effect converges to zero.
@@
Figure 2(b) shows the effect of MHSA knockout, which is similar to what we observed with MLP sublayer knockout. Specifically, knocking out early layers MHSA (0-3) leads to a drop in $\Delta F(s_{a})$ and increase in $\Delta L(s_{a})$. The most significant difference between $\Delta F(s_{a})$ and $\Delta L(s_{a})$ is in layer 1, where $\Delta F(s_{a})$ dropped by 0.30-0.30 and $\Delta L(s_{a})$ increased by $0.16$ ($p<0.05$). After layer 4, knocking out MHSA has little to no effect on interpretation, as the values converge toward zero.
@@
Together, the two interventions track the 1 idiom’s figurative interpretation retrieval: Early MHSA layers (0-3) gather and bind token-wise interactions, while early MLP layers (0-3) transform those bound representations into a figurative interpretation, while suppressing the literal one. This early layer idiom interpretation retrieval aligns with the findings of Haviv et al. (2023), who argue that idiomatic information is accessed in the initial layers of LLMs during inference. After layer 4, $\Delta F(s_{a})$ and $\Delta L(s_{a})$ stabilize to zero, indicating that later sublayers are not adding new information and that the representations of the interpretations at the idiom span tokens are no longer used by any subsequent tokens.
@@
Moreover, these components are specifically essential for idiom processing, not generally important for semantic interpretation of a sentence. Figure 2(c) shows that the drop from knocking out early MLP sublayers is significantly smaller in layer 1 for the two paraphrases ($\Delta L(s_{l})$, $\Delta F(s_{f})$) than for the idiomatic expression, ($\Delta F(s_{a})$), ($p<0.05$). While, at layer 0, the knockout effect is large for all expressions, as in this case, the MLP block is processing the semantics of the sentence. Figure 2(d) shows a similar effect for the early MHSA layers (0-2) where literal paraphrases are not strongly affected by knockout ($\Delta L(s_{l})$, $\Delta F(s_{f})$) compared to idiomatic expression ($\Delta F(s_{a})$), ($p<0.05$). The crucial role of figurative interpretation retrieval in early layers is consistent across different models (Appendix D).
4.2 Identifying attention heads specific to retrieving figurative interpretations